# IsoSEL: Isometric Structural Entropy Learning for Deep Graph Clustering in Hyperbolic Space

### v1

@@ IV Theory: Structural Information & Graph Clustering
In this section, we derive an equivalent form of differential structural information to reduce its complexity, and provide the new theoretical results on the advantages of structural information for graph clustering.
@@ IV-A Reducing Complexity
[…]
@@ IV-C Handling Imbalanced Clusters
Nodes within the cluster are typically densely connected while there are fewer links connecting the nodes of different clusters. The condition is easy to be satisfied in real graphs.
@@
V Isometric Graph Augmentation
The structural information solely handles the graph structure, while node attributes are often equally important to graph clustering. To fill this gap, we propose isometric augmentation which integrates attribute information into graph structure, generalizing existing methods from representation perspective.
V-A Isometry between Manifolds
We demonstrate that arbitrary embedding induces a Riemannian manifold of hypersurface, and the existence of this manifold is claimed in the following lemma.
Lemma V.1 (Existence).
Given a set of points $\{\boldsymbol{p}_{i}\}_{i=1,..,N}$ on $\mathbb{R}^{n+1}$ derived from certain embedding, there exists a hypersurface $\mathcal{M}$ such that $\boldsymbol{p}_{i}\in\mathcal{M}$ holds for all $i$.
Proof.
Refer to Appendix II, Lemma V.1. ∎
Now, we elaborate on the isometry between manifolds. Specifically, we leverage the notion of Gaussian curvature, which depicts the local geometry of the manifold. Considering a point $\boldsymbol{p}_{i}\in\mathcal{M}$ on the manifold, the Gaussian curvature at $\boldsymbol{p}$ is intrinsically defined by the rate at which the geodesics on the manifold of hypersurface, emanating from $\boldsymbol{p}$, spread apart and this curvature is invariant among coordinate systems. Two hypersurfaces are said to be isometric if their Gaussian curvature are preserved relative to each other according to Gauss’s Theorema Egregium [24]. We develop the geometric intuition: The isometry means that two hypersurfaces exhibit the same pattern on how it curves at each point.
To find an isometry between hypersurfaces, we present our theoretical result on Givens rotation matrix $\mathbf{G}_{\theta}$. To be specific, the element of $\mathbf{G}_{\theta}$ is defined as follows,
$\mathbf{G}_{\theta}[i,j]=\mathbf{I}[i,j]+\beta_{ii}(\cos\theta-1)+\beta_{ij}\sin\theta,\vspace{-0.03in}$ (22)
where $\beta_{ii}=e_{i}e_{i}^{\top}+e_{j}e_{j}^{\top}$, $\beta_{ij}=e_{i}e_{j}^{\top}+e_{j}e_{i}^{\top}$, $\mathbf{I}$ is the identity, and $e$’s denote the orthogonal bases of the hypersurface. The matrix $\mathbf{G}_{\theta}$ characterizes the rotation along the span of $e_{i}$ and $e_{j}$ with the rotation angle of $\theta$. We prove the following statement holds for arbitrary hypersurfaces. {mymath}
Theorem V.2 (Rotational Isometry).
Given two hypersurfaces $\mathcal{S}_{1},\mathcal{S}_{2}$ in $\mathbb{R}^{n+1}$, equipped with the usual inner product, the Givens rotation $\mathbf{G}_{\theta}$ can be an isometry, such that $\mathbf{G}_{\theta}:\mathcal{S}_{1}\rightarrow\mathcal{S}_{2}$, and vice versa.
Proof.
Refer to Appendix II, Theorem V.2. ∎
In other words, the Givens rotation instantiate an isometry between hypersurfaces in terms of Gauss’s Theorema Egregium.
V-B Augmentation with Rotational Isometry
Typical integration methods [25] create an augmented graph $\hat{G}(\hat{\mathbf{A}},\mathbf{X})$ that aggregates the adjacency matrix and the similarity matrix of node attributes, given a graph $G(\mathbf{A},\mathbf{X})$. The augmented $\hat{\mathbf{A}}$ is given by the adjacency matrix $\mathbf{A}$ and the similarity $\mathbf{S}$ derived from the attributes $\mathbf{X}$, e.g., the structure of K-nearest neighbor graph based on attribute similarity.
In IsoSEL, we propose a more generic approach (referred to as isometric augmentation) from the representation perspective. The key idea is to perform aggregation on the manifolds and derive the augmented structure with the notion of isometry. First, we induce the manifold of $G$ with message-passing embedding, taking the popular SGC [26] as an example. Hence two sets of embeddings are obtained from adjacency matrix and attribute similarity, formulated as $\mathbf{A}^{L}\mathbf{X}$ and $\mathbf{S}^{L}\mathbf{X}$, respectively. The message-passing is conducted $L$ times without loss of generality. Second, we select an aggregation function $f$ to construct the hypersurface whose existence is guaranteed given Lemma V.1. $f$ can be a weighted sum or product, and the corresponding embedding is written as $f(\mathbf{A},\mathbf{S},\mathbf{X};L)$. Next, we aim to derive an augmented graph whose induced hypersurface is isometric to that of $f(\mathbf{A},\mathbf{S},\mathbf{X};L)$ in terms of Gauss’s Theorema Egregium. However, specifying the relationship between the embedding of isometric hypersurfaces is challenging. Thanks to the discovery of Theorem V.2, the augmented $\hat{\mathbf{A}}$ can be formulated in the optimization problem as follows. {mymath}
Theorem V.3 (Optimization of Isometric Augmentation).
Given a graph with an adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ and a similarity matrix $\mathbf{S}\in\mathbb{R}^{N\times N}$ induced by the attributes $\mathbf{X}\in\mathbb{R}^{N\times d}$, the adjacency matrix of augmented graph $\hat{\mathbf{A}}\in\mathbb{R}^{N\times N}$ considering the isometry in the $L$ convolution is the closed-form minimizer of the following optimization
$\min\ \|f(\mathbf{A},\mathbf{S},\mathbf{X};L)\mathbf{G}^{\top}_{\theta}-\hat{\mathbf{A}}^{L}\mathbf{X}\|^{2},\vspace{-0.05in}$ (23)
subject to the symmetry matrix cone of $\hat{\mathbf{A}}$, where $\mathbf{G}_{\theta}\in\mathbb{R}^{d\times d}$ denotes the Givens rotation, and $f(\mathbf{A},\mathbf{S},\mathbf{X};L)\in\mathbb{R}^{N\times d}$ is the aggregator of structural and attribute representations.
Proof.
Refer to Appendix II, Theorem V.3. ∎
Closed-form Expression. The construction above enjoys the existence of closed-form solution. In particular, the augmented graph is expressed as follows,
$\hat{\mathbf{A}}=\mathbf{P}\operatorname{diag}(\lambda_{1}^{\frac{1}{L}},\cdots,\lambda_{d}^{\frac{1}{L}})\mathbf{P}^{\top},$ (24)
where $\mathbf{P}$ and $\lambda$’s are derived as the eigenvectors and eigenvalues of the following matrix
We have $\mathbf{B}=f(\mathbf{A},\mathbf{S},\mathbf{X};L)\mathbf{G}_{\theta}^{T}$, and $\mathbf{U}=\mathbf{X}\mathbf{X}^{T}+\omega\mathbf{I}$ with penalty coefficient $\omega$. Note that, the eigenvalue decomposition is well-established with efficient methods [27] while the augmentation is typically applied as a preprocessing. The advantage is that we improve the expressiveness of the polynomial over structural and attribute matrices by incorporating representation consideration, as shown in Ablation Study, while recovering previous methods [25] as special cases.
VI LSEnet: Learning Partitioning Tree in Hyperbolic Space
@@
We design a novel Lorentz Structural Entropy Neural Network (LSEnet), which aims to learn the optimal partitioning tree $\mathcal{T}_{\operatorname{net}}^{*}$ in the Lorentz model of hyperbolic space. LSEnet constructs the partitioning tree in a bottom-up fashion where we first embed leaf nodes of the tree, and then recursively learn the parent nodes, leveraging the clustering objective of differential structural information. In particular, we learn the partitioning trees from both the original and augmented structures, which serve as the different views in Lorentz tree contrastive learning (Sec. VII).
VI-A Why Hyperbolic Space?
@@
Theorem VI.1 (Tree $\mathcal{T}$ and Hyperbolic Space [28]).
@@
That is, hyperbolic space is well suited to embed the partitioning tree. Note that, Theorem VI.1 does not hold for Euclidean space or other Riemannian manifold.
VI-B Embedding Leaf Nodes
@@
$\operatorname{LLinear}(\boldsymbol{x})=\begin{bmatrix}\sqrt{\lVert h(\mathbf{W}\boldsymbol{x},\boldsymbol{v})\rVert^{2}-\frac{1}{\kappa}}\ \\ h(\mathbf{W}\boldsymbol{x},\boldsymbol{v})\ \end{bmatrix}\in\mathbb{L}^{\kappa,d_{\mathcal{T}}},\vspace{-0.03in}$ (26)
@@
$\omega_{ij}$ $=\operatorname{LAtt}(\mathbf{Q},\mathbf{K})=\frac{\exp(a_{ij})}{\sum_{(i,l)\in\mathcal{E}}\exp(a_{il})},\vspace{-0.03in}$ (28)
where $\mathbf{W}\in\mathbb{R}^{1\times 2d}$ is learnable parameters, [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] is concatenate operation, $\boldsymbol{q}$ and $\boldsymbol{k}$ are the query and key vector collected in row vector of $\mathbf{Q}$ and $\mathbf{K}$, respectively. The queries and keys are derived from node feature $\boldsymbol{x}$ via $\operatorname{LLinear}$ with different parameters. Third, weighted aggregation is considered as the arithmetic mean among manifold-valued vectors. Given a set of points $\{\boldsymbol{x}_{i}\}_{i=1,\cdots,N}$ in the Lorentz model $\mathbb{L}^{\kappa,d_{\mathcal{T}}}$,
$\operatorname{LAgg}(\boldsymbol{\omega},\mathbf{X})=\frac{1}{\sqrt{-\kappa}}\sum_{i=1}^{N}\frac{\omega_{i}}{\lvert\lVert\sum_{j=1}^{N}\omega_{j}\boldsymbol{x}_{j}\rVert_{\mathbb{L}}\rvert}\boldsymbol{x}_{i},\vspace{-0.03in}$ (29)
@@
VI-C Learning Parent Nodes
@@
[…]
$\mathbf{C}^{h}=\operatorname{LAtt}(\mathbf{Q},\mathbf{K})\sigma(\operatorname{MLP}(\mathbf{Z}^{h})),\vspace{-0.03in}$ (31)
@@
[…]
$\boldsymbol{z}^{h-1}_{j}=\arg_{\boldsymbol{z}^{h-1}_{j}}\min\sum\nolimits_{i=1}^{N}c_{ij}d^{2}_{\mathbb{L}}(\boldsymbol{z}^{h-1}_{j},\boldsymbol{z}^{h}_{i}).\vspace{-0.03in}$ (32)
@@
Theorem VI.2 (Geometric Centroid).
@@
$\boldsymbol{z}^{h-1}_{j}=\frac{1}{\sqrt{-\kappa}}\sum_{i=1}^{N}\frac{c_{ji}}{\lvert\lVert\sum_{l=1}^{n}c_{jl}\boldsymbol{z}^{h}_{l}\rVert_{\mathbb{L}}\rvert}\boldsymbol{z}^{h}_{i}\in\mathbb{L}^{\kappa,d_{\mathcal{T}}},\vspace{-0.03in}$ (34)
@@
Remark. In fact, the geometric centroid in Theorem VI.2 is also equivalent to the gyro-midpoint in Poincaré ball model of hyperbolic space, detailed in Appendix II, Theorem VI.2.
@@
$\mathbf{A}^{h-1}$ $=(\mathbf{C}^{h})^{\top}\mathbf{A}^{h}\mathbf{C}^{h}.\vspace{-0.03in}$ (35)
VI-D Hyperbolic Partitioning Tree
In the principle of structural information minimization, the optimal partitioning tree is constructed to describe the self-organization of the graph [16, 18]. In the continuous realm, LSEnet learns the optimal partitioning tree in hyperbolic space by minimizing the objective in Eq. (7). DSI is applied on all the level-wise assignment $\mathbf{C}$’s, which are parameterized by neural networks in Sec. VI-B and VI-C. We suggest to place the tree root at the origin of $\mathbb{L}^{\kappa,d_{\mathcal{T}}}$, so that the learnt tree enjoys the symmetry of Lorentz model. Consequently, the hyperbolic partitioning tree describes the graph’s self-organization and clustering structure in light of Theorem III.5.
VII Graph Clustering with
Lorentz Tree Contrastive Learning
In this section, we introduce the Lorentz Tree Contrastive Learning designed to contrast trees in the hyperbolic space, achieving two key objectives. First, it incorporates attribute information in a contrastive manner to overcome the limitations of structural entropy. Second, it refines the hyperbolic tree representations, thereby enhancing clustering assignment. The main components of Lorentz Tree Contrastive Learning are i) Lorentz projection head, ii) hyperbolic reweighed sampling and iii) the Lorentz contrastive loss.
VII-A Neural Transformation in Lorentz Group
The projection head for contrasting partitioning trees differs significantly from typical graph contrastive learning. The key distinctions are as follows. First, existing methods primarily explore similarity in Euclidean space, whereas the partitioning tree is aligned with hyperbolic space, presenting a manifold-preserving challenge. In other words, the projection head is required to connect two hyperbolic spaces. Second, in graph contrastive learning, the projection head operates on graph nodes (leafs) in graph contrastive learning, while in the tree contrastive learning, it manipulates tree nodes across all levels simultaneously. Notably, maintaining consistency between successor and ancestor nodes is crucial owing to the unity of the tree, as supported by empirical evidence in Sec. VIII-E. However, ensuring consistency in hyperbolic space remains a significant challenge.
Before elaborating on the proposed projection head, we revisit the notion of Lorentz group $G^{L}=\{L_{i}\}$. To be specific, Lorentz group is a set of Lorentz transformations $L$, and $L_{i}\circ L_{j}\in G^{L}$ is closed on the group for any $L_{i},L_{j}\in G^{L}$. For any $\boldsymbol{x}\in\mathbb{L}^{\kappa,d_{1}}$, the equality of $L(\boldsymbol{x})=\mathbf{L}\boldsymbol{x}\in\mathbb{L}^{\kappa,d_{2}}$ holds for the block diagonal $\mathbf{L}=\left[\begin{array}[]{cc}w&\boldsymbol{v}^{\top}\\ \boldsymbol{v}&\mathbf{W}\end{array}\right],$ provided that $\mathbf{W}$ characterizes a Lorentz boost or Lorentz rotation. The advantage of Lorentz transformation lies in its clean matrix-vector multiplication form. However, the condition on $\mathbf{W}$ makes the optimization problematic; for instance, Lorentz rotation is restricted to the orthogonal cone of $\mathbf{W}$.
To bridge this gap, we introduce a novel Lorentz-type group consisting of learnable transformations that (i) connect different hyperbolic spaces (manifold-preserving) and (ii) ensure consistency for tree contrastive learning. The transformation is written as $L(\boldsymbol{x})=\mathbf{L}\boldsymbol{x}\in\mathbb{L}^{\kappa,d_{2}}$ for arbitrary $\boldsymbol{x}\in\mathbb{L}^{\kappa,d_{1}}$, where $d_{1}$ and $d_{2}$ are not necessarily the same. $\mathbf{L}$ takes the form of $\mathbf{L}:=\begin{bmatrix}\mathbf{v}^{T}\\ \mathbf{W}\end{bmatrix}$ satisfying the equation of
$\mathbf{W}^{T}\mathbf{W}=\mathbf{v}\mathbf{v}^{T}+diag(-1,1,...,1)\vspace{-0.05in}$ (36)
with $\mathbf{W}\in\mathbb{R}^{d_{2}\times d_{1}}$, and $d_{2}>d_{1}$ in the projection head. Next, we show the theoretical results on our construction. {mymath}
Theorem VII.1 (Manifold-preserving).
Given the construction of $\mathbf{L}$ in Eq. (36) and scalars $a,b$, the following properties holds for any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{L}^{d_{1}}$.
@@ VII-A Neural Transformation in Lorentz Group
where $\langle\boldsymbol{x},\boldsymbol{y}\rangle_{\mathbb{L}}$ denotes the Minkowski inner product in Eq. (6).
@@
Refer to Appendix II, Theorem VII.1. ∎
Note that, the manifold-preserving property is given by the claim (2) in the theorem above. It is easy to verify that $G$ is closed on itself, i.e., $L_{i}\circ L_{j}\in G^{L}$, $\forall L_{i},L_{j}\in G^{L}$. These properties also induce the following result. {mymath}
Theorem VII.2 (Consistence Insurance).
Given a set of points $\{\mathbf{z}_{i}\}$ in the Lorentz model of hyperbolic space, $\mathbf{z}_{i}\in\mathbb{L}^{\kappa,d_{1}}$ and their midpoint $\mathbf{z}\in\mathbb{L}^{\kappa,d_{1}}$ is constructed by Eq. (29), then $L(\mathbf{z})\in\mathbb{L}^{\kappa,d_{2}}$ is the midpoint of $\{L(\mathbf{z}_{i})\}$, $L(\mathbf{z}_{i})\in\mathbb{L}^{\kappa,d_{2}}$, if $L$ is a transformation in the Lorentz group as stated in Theorem VII.1.
@@
Refer to Appendix II, Theorem VII.2. ∎
@@
Closed-form Expression. Another advantage is that Eq. (36) induces the closed-form expression. Specifically, we parameterize $\mathbf{L}$ on $\mathbf{v}$, and the corresponding $\mathbf{W}$ is derived as follows,
$\mathbf{W}=\begin{bmatrix}\mathbf{P}diag(\sqrt{\lambda_{1}},..,\sqrt{\lambda_{d+1}})\mathbf{P}^{T}\\ \mathbf{0}\end{bmatrix},\vspace{-0.05in}$ (37)
where $\mathbf{P}$ and $\lambda$’s are obtained from eigenvalue decomposition of $\mathbf{V}=\mathbf{v}\mathbf{v}^{T}+diag(-1,1,...,1)$ which takes the form of
$\mathbf{V}$ $=\mathbf{P}diag(\lambda_{1},..,\lambda_{d+1})\mathbf{P}^{T},\vspace{-0.05in}$ (38)
with eigenvalues $\lambda_{1},..,\lambda_{d+1}$ sorted in descending order. $\mathbf{0}$ is the padding matrix of zeros for dimensionality consistence.

### v2

@@ IV Theory: Structural Information & Graph Clustering
In this section, we derive a more efficient formulation of differential structural information, achieving linear complexity, and provide the new theoretical results on the advantages of structural information for graph clustering.
@@ IV-A Reducing Complexity
where $G^{h-1}$ is induced by the assignment matrix $\mathbf{S}$ in Eq. (8) at the (h1)(h-1) level, and $d_{i}^{h}$ is the degree of $i$-th node in graph $G^{h}$. The equivalence and variables in Eq. (17) are stated in the following theorem. {mymath}
@@ IV-C Handling Imbalanced Clusters
Nodes within the cluster are typically densely connected while there are fewer links connecting the nodes of different clusters. The condition is easy to satisfy in real graphs.
@@
V LSEnet: Learning Partitioning Tree in Hyperbolic Space
@@
We design a novel Lorentz Structural Entropy Neural Network (LSEnet), which aims to learn the optimal partitioning tree $\mathcal{T}_{\operatorname{net}}^{*}$ in the Lorentz model of hyperbolic space. LSEnet constructs the partitioning tree in a bottom-up fashion where we first embed leaf nodes of the tree, and then recursively learn the parent nodes, leveraging the clustering objective of differential structural information. In particular, we learn the partitioning trees from both the original and augmented structures, which serve as the different views in Lorentz tree contrastive learning (Sec. VI).
V-A Why Hyperbolic Space?
@@
Theorem V.1 (Tree $\mathcal{T}$ and Hyperbolic Space [48]).
@@
That is, hyperbolic space is well suited to embed the partitioning tree. Note that, Theorem V.1 does not hold for Euclidean space or other Riemannian manifold.
V-B Embedding Leaf Nodes
@@
$\operatorname{LLinear}(\boldsymbol{x})=\begin{bmatrix}\sqrt{\lVert h(\mathbf{W}\boldsymbol{x},\boldsymbol{v})\rVert^{2}-\frac{1}{\kappa}}\ \\ h(\mathbf{W}\boldsymbol{x},\boldsymbol{v})\ \end{bmatrix}\in\mathbb{L}^{\kappa,d_{\mathcal{T}}},\vskip-2.168pt$ (22)
@@
$\omega_{ij}$ $=\operatorname{LAtt}(\mathbf{Q},\mathbf{K})=\frac{\exp(a_{ij})}{\sum_{(i,l)\in\mathcal{E}}\exp(a_{il})},\vskip-2.168pt$ (24)
where $\mathbf{W}\in\mathbb{R}^{1\times 2d}$ is learnable parameters, $\cdot$$\cdot$] is concatenate operation, $\boldsymbol{q}$ and $\boldsymbol{k}$ are the query and key vector collected in row vector of $\mathbf{Q}$ and $\mathbf{K}$, respectively. The queries and keys are derived from node feature $\boldsymbol{x}$ via $\operatorname{LLinear}$ with different parameters. Third, weighted aggregation is considered as the arithmetic mean among manifold-valued vectors. Given a set of points $\{\boldsymbol{x}_{i}\}_{i=1,\cdots,N}$ in the Lorentz model $\mathbb{L}^{\kappa,d_{\mathcal{T}}}$,
$\operatorname{LAgg}(\boldsymbol{\omega},\mathbf{X})=\frac{1}{\sqrt{-\kappa}}\sum_{i=1}^{N}\frac{\omega_{i}}{\lvert\lVert\sum_{j=1}^{N}\omega_{j}\boldsymbol{x}_{j}\rVert_{\mathbb{L}}\rvert}\boldsymbol{x}_{i},\vskip-2.168pt$ (25)
@@
$\operatorname{LConv}(\mathbf{X}|\mathbf{A})=\operatorname{LAgg}(\operatorname{LAtt}(\mathbf{Q},\mathbf{K})\odot\mathbf{A},\operatorname{LLinear}(\mathbf{X})),$
@@
V-C Learning Parent Nodes
@@
Without loss of generality, we consider the assignment between $h$-th and (h1)(h-1)-th levels given node embeddings $\boldsymbol{z}^{h}\in\mathbb{L}^{\kappa,d_{\mathcal{T}}}$ at $h$-th level. Recalling Definition III.1, the $i$-th row of assignment $\mathbf{C}^{h}\in\mathbb{R}^{N_{h}\times N_{h-1}}$ describes the belonging of $i$-th node at $h$-th to the parent nodes at (h1)(h-1)-th. We design a Lorentz assigner following the intuition that neighborhood nodes in the graph tend to have similar assignments. Concretely, we leverage Multilayer Perceptron $\operatorname{MLP}$ to learn the assignment from embeddings, and meanwhile the similarity is parameterized by $\operatorname{LAtt}$ defined in Eq. 24. Thus, the Lorentz assigner is formulated as follows,
$\mathbf{C}^{h}=\operatorname{LAtt}(\mathbf{Q},\mathbf{K})\sigma(\operatorname{MLP}(\mathbf{Z}^{h})),\vskip-2.168pt$ (27)
@@
The remaining task is to infer the node embeddings $\boldsymbol{z}^{h-1}\in\mathbb{L}^{\kappa,d_{\mathcal{T}}}$ at (h1)(h-1)-th level. As reported in [7], a parent node is located at the point that has the shortest path to all the child nodes at the $h$-th level, correspondingly, the parent node is the Fréchet mean of the child nodes in the manifold. The challenge here is that Fréchet mean regarding the canonical distance has no closed-form solution [40]. Alternatively, we consider the geometric centroid for the squared Lorentz distance, where the weights are given by the soft assignment $\mathbf{C}^{h}$,
$\boldsymbol{z}^{h-1}_{j}=\arg_{\boldsymbol{z}^{h-1}_{j}}\min\sum\nolimits_{i=1}^{N}c_{ij}d^{2}_{\mathbb{L}}(\boldsymbol{z}^{h-1}_{j},\boldsymbol{z}^{h}_{i}).\vskip-2.168pt$ (28)
@@
Theorem V.2 (Geometric Centroid).
@@
$\boldsymbol{z}^{h-1}_{j}=\frac{1}{\sqrt{-\kappa}}\sum_{i=1}^{N}\frac{c_{ji}}{\lvert\lVert\sum_{l=1}^{n}c_{jl}\boldsymbol{z}^{h}_{l}\rVert_{\mathbb{L}}\rvert}\boldsymbol{z}^{h}_{i}\in\mathbb{L}^{\kappa,d_{\mathcal{T}}},\vskip-2.168pt$ (30)
@@
Remark. In fact, the geometric centroid in Theorem V.2 is also equivalent to the gyro-midpoint in Poincaré ball model of hyperbolic space, detailed in Appendix II, Theorem VI.2. Note that attentive aggregation for parent node embedding is uni-directional, different from that in typical graph neural nets.
According to Eq. (27), it requires the graph structure $\mathbf{A}^{h-1}$ among the parent nodes at (h1)(h-1)-th level to derive the assignment. We give the adjacency matrix of the (h1)(h-1)-th level graph as follows,
$\mathbf{A}^{h-1}$ $=(\mathbf{C}^{h})^{\top}\mathbf{A}^{h}\mathbf{C}^{h}.\vskip-2.168pt$ (31)
V-D Hyperbolic Partitioning Tree
In the principle of structural information minimization, the optimal partitioning tree is constructed to describe the self-organization of the graph [27, 63]. In the continuous realm, LSEnet learns the optimal partitioning tree in hyperbolic space by minimizing the objective in Eq. (7). DSI is applied on all the level-wise assignment $\mathbf{C}$’s, which are parameterized by neural networks in Sec. V-B and V-C. We suggest to place the tree root at the origin of $\mathbb{L}^{\kappa,d_{\mathcal{T}}}$, so that the learnt tree enjoys the symmetry of Lorentz model. Consequently, the hyperbolic partitioning tree describes the graph’s self-organization and clustering structure in light of Theorem III.5.
VI Lorentz Tree Contrastive Learning
The partitioning tree above resides in the hyperbolic space, and thus we propose to refine its hyperbolic representation through Lorentz tree contrastive learning, enhancing the graph semantics from a representational perspective. In this contrastive learning framework, we develop a novel neural Lorentz boost for consistency insurance and a tree contrastive loss considering the tree branches in light of clustering assignment.
VI-A Neural Lorentz Boost
Different from graph contrastive learning, the projection head in our case further requires the consistency insurance—maintaining the parent hierarchies of the partitioning tree as a unity. That is, the level-wise child-parent assignment cannot be disrupted by this projection head.
To achieve this, we leverage the Lorentz group $G^{L}=\{L_{i}\}$, whose elements are isometries over Lorentz model. For any $\boldsymbol{x}\in\mathbb{L}^{\kappa,d}$, a Lorentz transformation can be written as follows,
@@ VI-A Neural Lorentz Boost
where $\mathbf{W}\in\mathbb{R}^{d\times d}$ is a matrix, $w$ is a scaling scalar, and $\boldsymbol{v}\in\mathbb{R}^{d}$ is a boost vector. Note that Lorentz rotations act tangentially to the radial direction. They rotate the partitioning tree and do not affect representation refinement due to rotational symmetry. Thus, we consider the Lorentz boosts, which serve as a linear transformation to operate on the partitioning tree in the Lorentz model of hyperbolic space. We derive a closed-form Lorentz boost parameterized by $\boldsymbol{\beta}\in\mathbb{R}^{d}$, maintaining the consistency insurance in partitioning tree contrasting (as shown in Theorems VI.1 and VI.2).
{mymath}
Theorem VI.1 (Closed-from Lorentz Boost).
Given the construction in Eq. (32) and a parameter vector $\boldsymbol{\beta}\in\mathbb{R}^{d}$, Lorentz boosts have a closed form with components given as
$w$ $=\frac{1}{\sqrt{1-\|\boldsymbol{\beta}\|^{2}_{\mathbb{L}}}}$ (33)
$\boldsymbol{v}$ $=-w\cdot\boldsymbol{\beta}$ (34)
$\mathbf{W}$ $=\frac{(w-1)}{\|\boldsymbol{\beta}\|^{2}_{\mathbb{L}}}\cdot\boldsymbol{\beta}\boldsymbol{\beta}^{\top}+\mathbf{I}_{d}$ (35)
Proof.
Refer to Appendix II.G, Theorem VI.1. ∎
{mymath}
Theorem VI.2 (Manifold Preserving).
Given the construction of $\mathbf{L}$ in Theorem VI.1, and scalars $a,b$, the following properties holds for any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{L}^{d_{1}}$.
@@
where $\langle\boldsymbol{x},\boldsymbol{y}\rangle_{\mathbb{L}}$ denotes the Minkowski inner product.
@@
Refer to Appendix II.G, Theorem VI.2. ∎
Note that, the manifold-preserving property is given by claim (2) in the theorem above. It is easy to verify that $G$ is closed on itself, i.e., $L_{i}\circ L_{j}\in G^{L}$, $\forall L_{i},L_{j}\in G^{L}$. These properties also induce the following result. {mymath}
Theorem VI.3 (Consistence Insurance).
Given a set of points $\{\boldsymbol{z}_{i}\}$ in the Lorentz model of hyperbolic space, $\boldsymbol{z}_{i}\in\mathbb{L}^{\kappa,d_{1}}$ and their midpoint $\boldsymbol{z}\in\mathbb{L}^{\kappa,d_{1}}$ is constructed by Eq. (25), then $L(\boldsymbol{z})\in\mathbb{L}^{\kappa,d_{2}}$ is the midpoint of $\{L(\boldsymbol{z}_{i})\}$, $L(\boldsymbol{z}_{i})\in\mathbb{L}^{\kappa,d_{2}}$, if $L$ is a transformation in the Lorentz group as stated in Theorem VI.2.
@@
Refer to Appendix II.H, Theorem VI.3. ∎